# Bayesian Intrinsic Groupwise Image Registration: Unsupervised Disentanglement of Anatomy and Geometry

### v1

@@ 4 An Interpretable Registration Model via Bayesian Disentanglement Learning
Fig. 3 presents a schematic illustration of the proposed interpretable registration network, specially designed based on the optimization procedure for Bayesian inference outlined in the previous section. The encoders first disentangle the input multi-modal images into structural representation maps with categorical distribution. Then an average representation of the common anatomy is computed within the latent space, so that spatial correspondence between the observed images and the common anatomy are predicted from these structural representations. The disentanglement is learned by means of inversely simulating the image generative process $p(\bm{u}\mid\bm{z},\bm{v};\bm{\theta})$ with decoders, i.e. reconstructing the original images based on the common anatomy representation. Fig. 4 depicts the entire network architecture, composed of the encoders, the registration modules and the decoders. The following subsections will detail the construction of these modules.
@@ 4.1 Inference of Structural Representations
In the inference steps #1 (ref. Fig. 0(b)), given the original misaligned image group $\bm{u}$, the encoders first estimate the multi-level categorical distributions $\widetilde{q}_{j}(\bm{z}^{l}\mid u_{j};{\bm{\psi}_{j}})$ for each original image $u_{j}$, where $l$ indicates the spatial resolution of the representation. Note that $\widetilde{q}_{j}(\bm{z}^{l}\mid u_{j};{\bm{\psi}_{j}})\triangleq\operatorname{Cat}(\bm{z};\widetilde{\bm{p}}_{j,1}^{l},\dots,\widetilde{\bm{p}}_{j,K}^{l})$ differs from $q_{j}^{\diamond}(\bm{z}^{l}\mid u_{j},\bm{v}_{j};{\bm{\psi}_{j}})=\operatorname{Cat}(\bm{z};\bm{\pi}_{j,1}^{\diamond\,l},\dots,\bm{\pi}_{j,K}^{\diamond\,l})$ regarding the network input. The former takes $\bm{u}=\{u_{j}\}_{j=1}^{N}$ as input, while the latter takes $\bm{u}\circ\bm{\phi}=\{u_{j}\circ\phi_{j}\}_{j=1}^{N}$. The architecture of the encoders is adapted from the Attention U-Net [82], which enhances the vanilla U-Net [83] with additive attention gates. Moreover, domain-specific batch normalization [84] is utilized in the contracting path of the encoders to cancel out modality-specific appearance variations, while all convolutional layers are shared across modalities, embedding multi-modal images into modality-invariant representations.
@@ 4.2 Inference of Velocity Fields
Based on the structural representation maps extracted from the original images by the encoders, the registration modules seek to estimate the transformations that spatially register the image group. This is realized in a coarse-to-fine approach, starting from prediction of the lowest-resolution velocity fields, (ref. the red arrows in Fig. 0(b)).
@@
As the structural representation maps $\widetilde{\bm{p}}_{j}^{l}\circ\phi_{j}^{<l}$ encode the structural features of the partially registered input images, we can in fact compute the mean $\bm{\mu}_{\bm{v},j}^{l}$ of the variational posterior $q(\bm{v}_{j}^{l}\mid\bm{u},\bm{v}^{<l};{\bm{\psi}})$ directly from them based on the Demons algorithm [85, 86, 87, 88]. Particularly, for each grid location $\bm{\omega}$ at level $l$, we define $\bm{\mu}_{\bm{v},j}^{l}$ by
@@
$\bm{\varphi}_{j,\bm{\omega}}^{l}({\bm{u}})\triangleq\widetilde{\bm{p}}^{l*}(\bm{\omega})-\widetilde{\bm{p}}_{j}^{l}\circ\phi_{j}^{<l}\circ(\operatorname{id}+\bm{u})(\bm{\omega}),$
@@
is the sample variance of the feature difference norm. Detailed derivation can be found in Appendix C, where the Demons algorithm is established from a mathematically interpretable optimization framework. The variance $\bm{\Sigma}_{\bm{v},j}^{l}$ of the variational posterior is predicted by a convolutional block from the concatenation of probability maps $[\widetilde{\bm{p}}_{j}^{l}\circ\phi_{j}^{<l};\widetilde{\bm{p}}^{l*}]$, whereby the network parameters are shared among different modalities.
In addition, to avoid the degeneracy that deforms the observed images to an arbitrary coordinate space, we constrain the sum of the final velocity fields to be zero by subtracting their average [17], i.e.
$\bm{v}_{j}^{+}\leftarrow\bm{v}_{j}^{+}-\frac{1}{N}\sum_{j=1}^{N}\bm{v}_{j}^{+}.$ (21)
@@ 4.3 Bayesian Disentangled Representation Learning
Given the perceptual modules capturing intrinsic structural representations from multi-modal images and their spatial correspondence, we are going to discuss how such mapping functions are learned in the first place. This is achieved by optimizing the network parameters $\bm{\psi}$ and $\bm{\theta}$ in a closed-loop self-reconstruction process that maximizes the ELBO. The single-view structural distributions for estimating the common anatomy are defined by
@@
Based on the common structural representations sampled from the geometric mean $q^{*}(\bm{z}^{l}\mid\bm{u},\bm{v};{\bm{\psi}})$, the decoders reconstruct the original images in two steps (ref. Fig. 0(d)). The images corresponding to the common anatomy $\widetilde{x}_{j}=f(\bm{z};\bm{\theta}_{j})$ are decoded first. Then, the estimation of the original images $\widehat{\bm{u}}=\{\widehat{x}_{j}\}_{j=1}^{N}$ are obtained by inverse warping, i.e. $\widehat{x}_{j}=\widetilde{x}_{j}\circ\phi_{j}^{-1}$ with $\phi_{j}^{-1}=\exp(-\bm{v}_{j}^{+})$, simulating Eq. 3.
The network architecture of the decoders is given by multi-level convolutional blocks and linear upsampling. At level $l$, the output feature maps from upsampling are multiplied by the sampled categorical variables $\bm{z}^{l}$. Besides, all convolutional layers are assigned with kernel size of 1 isotropically, to suppress spatial correlation and to form spatial equivariance. Thus, the decoded image $\widetilde{x}_{j}$ is in the same spatial orientation as the common anatomy $\bm{z}$, encouraging the registration module to identify the correct velocity fields $\bm{v}_{j}^{+}$ for reconstructing $\widehat{x}_{j}$. This connection between the spatial equivariance and the identifiability of registration is formally established in Appendix E. Furthermore, the batch normalization layers in the decoders are set as the inverse of the counterpart BN functions along the contracting path of the encoders, i.e.
@@
where for each image index $j$ and its corresponding modality $m$, $\bm{c}_{j}^{\text{in}},\bm{c}_{j}^{\text{out}}\in\mathds{R}^{B_{m}\times C\times H\times W\times D}$ are the input and output feature maps, $\bm{\mu}_{m},\bm{\sigma}_{m}\in\mathds{R}^{C}$ are the batch statistics, and $\bm{\beta}_{m},\bm{\gamma}_{m}\in\mathds{R}^{C}$ are the affine parameters. The arithmetic is computed element-wise. In fact, such preference for invertible and spatially decomposable decoders may promote learning of the true disentangled representations [10, 89]. An illustration of the encoder and decoder architecture can be found in Appendix D.
Finally, the expectation $\mathbb{E}_{q(\bm{z},\bm{v}\mid\bm{u};\bm{\psi})}[\log p(\bm{u}\mid\bm{z},\bm{v};{\bm{\theta}})]$ is approximated by Monte Carlo with ancestral sampling, where the likelihood is modelled by a Laplace distribution, i.e.,
$p(\bm{u}\mid\bm{z},\bm{v};\bm{\theta})$ $=\prod_{j=1}^{N}p(u_{j}\mid\bm{z},\bm{v}_{j};\bm{\theta}_{j})$ (22)
$\propto\prod_{j=1}^{N}\prod_{\bm{\omega}\in\Omega_{j}}\exp\left\{-\frac{\absolutevalue{u_{j}(\bm{\omega})-f(\bm{z};\bm{\theta}_{j})\circ\phi_{j}^{-1}(\bm{\omega})}}{b}\right\},$
@@
The proposed framework also encourages learning of disentangled representations formally defined in [8], which formulates disentangled representations as a reflection on the decomposition of certain symmetry groups acting on the world states (or observations). This explains why we choose to decompose the latent variables of our model into the common anatomy and the corresponding spatial diffeomorphisms: because they reflect two separate symmetry structures of the observed images, i.e., the ontological transformation changing the underlying anatomy, and the diffeomorphic transformation characterizing the geometric variation. More thorough discussion on disentangled representations and its identifiability can be found in Appendix E.
@@ 4.4 Learning, Inference and Scalability
Learning of our model is performed in an end-to-end fashion. All model parameters, including the auto-encoders and the registration modules, are trained simultaneously by stochastic gradient ascent to maximize the ELBO. At the inference stage, the transformations that register the image group is given by $\widehat{\bm{\phi}}=\{\widehat{\phi}_{j}\}_{j=1}^{N}$, where $\widehat{\phi}_{j}=\exp(\widehat{\bm{\mu}}_{j}^{+})$ and $\widehat{\bm{\mu}}_{j}^{+}$ is aggregated from the means of the velocity distributions at all levels predicted by the registration modules.
Remarkably, the proposed disentanglement of anatomy and geometry endows our model with powerful scalability. Since geometric and arithmetic averaging can be performed on any number (greater than 1) of representations, our model is capable of processing image groups of arbitrary sizes during training and test. Let the number of images in an input group be $N^{\prime}=\sum_{m=1}^{M}{N_{m}}$, where $N_{m}$ is the number of images of modality $m$. Each image of modality $m$ is processed by the encoder for the corresponding modality to extract structural representations, and then the geometric and arithmetic means are calculated on the representations of all images to estimate the KL divergence and predict deformations for each image. For decoding, the reconstruction from the decoder of modality $m$ is inversely warped for $N_{m}$ times by the deformations for the $N_{m}$ images of modality $m$ to reconstruct original images. Thus, our model allows $N_{m}$’s to be different or even 0, and only requires $N^{\prime}\geq 2$ (for calculating averages) and at least two modalities are available in each group, i.e., $\sum_{m}\text{1}(N_{m}>0)\geq 2$ (for effectively learning modality-invariant representations).
@@
  • During evaluation, our model can register groups with arbitrary sizes, enabling large-scale and variable-size multi-modal groupwise registration, as explored in Section 5.6. In contrast, conventional iterative methods optimize similarity metrics with computational complexities that escalate rapidly with group size. Deep learning baselines, on the other hand, often have fixed channel numbers to take as input the concatenation of images [90], necessitating groups to be of uniform size. These limitations significantly restrict their applicability.
@@ 5 Experiments and Results
We evaluated our framework on a total of four publicly available datasets, including BraTS-2021 [91, 92, 93], MS-CMRSeg [94], Learn2Reg Abdominal MR-CT [95, 96, 97], and the OASIS dataset [98]. The experiments were implemented in PyTorch [99] and conducted on $\text{NVIDIA}^{\circledR}$ $\text{RTX}^{\textnormal{TM}}$ 3090 GPUs. The experimental materials, baselines, evaluation metrics and results are detailed as follows.
@@ 5.2 Compared Methods
We compared two types of baselines with our models in the experiments. The first type features state-of-the-art iterative methods for multi-modal groupwise registration, including the accumulated pairwise estimates (APE) [4], conditional template entropy (CTE) [5], and $\mathcal{X}$-CoReg [6]. These methods do not need training, but they can be prone to image artifacts, and can hardly register large image groups due to excessive computational burdens. The second type of baselines extends the first type, aiming to learn the groupwise spatial correspondence via a neural network using the aforementioned similarity measures. To promote fair comparison, we used the Attention U-Net [82] for the network backbone, the same architecture as the encoder in our proposed framework. However, note that these learning-based baselines can only be trained to register images of a fixed group size, and therefore cannot be applied to data of variable group sizes in the test stage.
@@
  • Ours-CN: with complete learning and convolutional network based registration modules.
  • Ours-PD: with partial learning and Demons based registration modules.
  • Ours-CD: with complete learning and Demons based registration modules.
Here, “C” and “P” signify Complete and Partial learning strategies, respectively; “N” and “D” indicate Network or Demons based registration modules, respectively.
@@ 5.3 Implementation Details
During preprocessing, the intensity range of each image was linearly normalized to $[0,1]$. For the cardiac and abdominal images, they were further multiplied by the ROI mask to confine the information used by the network to predict registration. The encoder has an Attention U-Net architecture with $L=5$ levels, and the $l$-th level consists of two Conv-BN-LeakyReLU blocks with kernel size 3 and $16\times 2^{L-l}$ channels. The output of each convolutional block in the upsampling path is passed to a Conv layer to produce logits of the categorical structural distribution, with $8\times 2^{L-l}$ channels. Samples from the geometric mean structural distribution are passed to the decoder to reconstruct the original images. The decoder is composed of Conv-BN-LeakyReLU-Upsample blocks with kernel size 1 and $8\times 2^{L-l}$ channels.
To facilitate learning, different weights were used for the reconstruction loss, intrinsic structural distance, and registration regularization, corresponding to a likelihood-tempered ELBO [103]. Table II presents the balancing weights of the loss terms for different training datasets. We also added a local normalized cross-correlation (LNCC) term [104] to improve reconstruction on brain images, with its weight marked by values in the parentheses. The hyperparameter for spatial regularization was set to $\lambda=10$ as in [46], and the maximum magnitude of the velocity computed by the Demons algorithm was set to $\alpha_{0}^{l}=10\times 2^{l-L}$ at each level. The number of samples for the GR estimator was set to $S=3$ in all experiments. The network optimization was performed by stochastic gradient descent using the Adam optimizer [105]. Besides, we used learning rate of $1\times 10^{-3}$, and batch size of 20 or 2 for the MS-CMRSeg or the other datasets, respectively.
TABLE II: This table presents the respective weights of the reconstruction loss, intrinsic structural distance and registration regularization for different experimental setups.
Method
BraTS
MS-CMR
Learn2Reg
OASIS
Ours-PN
120 (60), 75, 15
120, 160, 10
N/A
N/A
Ours-CN
120 (60), 75, 20
120, 160, 10
120, 80, 4
120 (20), 40, 4
Ours-PD
120 (60), 120, 8
120, 160, 10
N/A
N/A
Ours-CD
120 (60), 60, 8
120, 160, 10
120, 80, 1
120 (20), 20, 0.1

### v2

@@ 4 An Interpretable Registration Architecture via Bayesian Disentanglement Learning
Fig. 3 presents a schematic overview of the proposed interpretable registration architecture, specially designed based on the optimisation procedure for Bayesian inference outlined in the previous sections and Fig. 1. The encoders first disentangle the input multi-modal images into structural representation maps with categorical distribution. Then an average representation of the common anatomy is computed within the latent space, so that spatial correspondence between the observed images and the common anatomy is predicted from these structural representations using the Diffeomorphic Demons algorithm. The disentanglement of geometric and anatomical variations is facilitated by inversely simulating the image generative process with spatially equivariant decoders, i.e., reconstructing the original images from the common anatomy representation. Fig. 4 depicts the entire network architecture, composed of the encoders, the registration modules and the decoders. The following subsections will detail the construction of these modules.
@@ 4.1 Inference of Structural Representations
In the inference steps #1 (ref. Fig. 1(b)), given the original misaligned image group $\bm{u}$, the encoders first estimate the multi-level categorical distributions $\widetilde{q}_{j}(\bm{z}^{l}\mid u_{j};{\bm{\psi}_{j}})$ for each original image $u_{j}$, where $l$ indicates the spatial resolution of the representation. Note that $\widetilde{q}_{j}(\bm{z}^{l}\mid u_{j};{\bm{\psi}_{j}})\triangleq\operatorname{Cat}(\bm{z};\widetilde{\bm{p}}_{j,1}^{l},\dots,\widetilde{\bm{p}}_{j,K}^{l})$ differs from $q_{j}^{\diamond}(\bm{z}^{l}\mid u_{j},\bm{v}_{j};{\bm{\psi}_{j}})=\operatorname{Cat}(\bm{z};\bm{\pi}_{j,1}^{\diamond\,l},\dots,\bm{\pi}_{j,K}^{\diamond\,l})$ regarding the network input. The former takes $\bm{u}=\{u_{j}\}_{j=1}^{N}$ as input, while the latter takes $\bm{u}\circ\bm{\phi}=\{u_{j}\circ\phi_{j}\}_{j=1}^{N}$. The architecture of the encoders is adapted from the Attention U-Net [83], which enhances the vanilla U-Net [84] with additive attention gates. Moreover, domain-specific batch normalisation [85] is utilised in the contracting path of the encoders to cancel out modality-specific appearance variations, while all convolutional layers are shared across modalities, embedding multi-modal images into modality-invariant representations.
@@ 4.2 Inference of Velocity Fields
Based on the structural representation maps extracted from the original images by the encoders, the registration modules seek to estimate the transformations that spatially register the image group. This is realised in a coarse-to-fine approach, starting from prediction of the lowest-resolution velocity fields (ref. the red arrows in Fig. 1(b)).
@@
As the structural representation maps $\widetilde{\bm{p}}_{j}^{l}\circ\phi_{j}^{<l}$ encode the structural features of the partially registered input images, we can in fact compute the mean $\bm{\mu}_{\bm{v},j}^{l}$ of the variational posterior $q(\bm{v}_{j}^{l}\mid\bm{u},\bm{v}^{<l};{\bm{\psi}})$ directly from them based on the Diffeomorphic Demons algorithm [86, 87]. Particularly, for each grid location $\bm{\omega}$ at level $l$, we define $\bm{\mu}_{\bm{v},j}^{l}$ by
@@
$\bm{\varphi}_{j,\bm{\omega}}^{l}({\bm{u}})\triangleq\widetilde{\bm{p}}^{l*}(\bm{\omega})-\widetilde{\bm{p}}_{j}^{l}\circ\phi_{j}^{<l}\circ(\operatorname{id}+\bm{u})(\bm{\omega})$
@@
is the sample variance of the feature difference norm. Detailed derivation can be found in Section C of the Supplementary Material, where the Demons algorithm is established from a mathematically interpretable optimisation framework. The variance $\bm{\Sigma}_{\bm{v},j}^{l}$ of the variational posterior is predicted by a convolutional block from the concatenation of probability maps $[\widetilde{\bm{p}}_{j}^{l}\circ\phi_{j}^{<l};\widetilde{\bm{p}}^{l*}]$, where the network parameters are shared among different modalities.
In addition, to avoid the degeneracy that deforms the observed images to an arbitrary coordinate space, we constrain the sum of the final velocity fields to be zero by subtracting their average [17], i.e., $\bm{v}_{j}^{+}\leftarrow\bm{v}_{j}^{+}-\frac{1}{N}\sum_{j=1}^{N}\bm{v}_{j}^{+}.$
@@ 4.3 Bayesian Disentangled Representation Learning
Given the perceptual modules capturing intrinsic structural representations from multi-modal images and their spatial correspondence, we are going to discuss how such mapping functions are learnt in the first place. This is achieved by optimising the network parameters $\bm{\psi}$ and $\bm{\theta}$ in a closed-loop self-reconstruction process that maximises ELBO. The single-view structural distributions for estimating the common anatomy are defined by
@@
Based on the common structural representations sampled from the geometric mean $q^{*}(\bm{z}^{l}\mid\bm{u},\bm{v};{\bm{\psi}})$, the decoders reconstruct the original images in two steps (ref. Fig. 1(d)). The images corresponding to the common anatomy $\widetilde{u}_{j}=f(\bm{z};\bm{\theta}_{j})$ are decoded first. Then, estimation of the original images $\widehat{\bm{u}}=\{\widehat{u}_{j}\}_{j=1}^{N}$ is obtained by inverse warping, i.e., $\widehat{u}_{j}=\widetilde{u}_{j}\circ\phi_{j}^{-1}$ with $\phi_{j}^{-1}=\exp(-\bm{v}_{j}^{+})$, simulating Eq. 3.
The network architecture of the decoders is given by multi-level convolutional blocks and linear upsampling. At level $l$, the output feature maps from upsampling are multiplied by the sampled categorical variables $\bm{z}^{l}$. Besides, all convolutional layers are assigned with a kernel size of 1 isotropically, to suppress spatial correlation and to form spatial equivariance. Thus, the decoded image $\widetilde{u}_{j}$ is in the same spatial orientation as the common anatomy $\bm{z}$, encouraging the registration module to identify the correct velocity fields $\bm{v}_{j}^{+}$ for reconstructing $\widehat{u}_{j}$. This connection between spatial equivariance and registration identifiability is formally established in Section E of the Supplementary Material. Furthermore, the batch normalization layers in the decoders are set as the inverse of the counterpart BN functions along the contracting path of the encoders, i.e.,
@@
where for each image index $j$ and its corresponding modality $m$, $\bm{c}_{j}^{\text{in}},\bm{c}_{j}^{\text{out}}\in\mathds{R}^{B_{m}\times C\times H\times W\times D}$ are the input and output feature maps, $\bm{\mu}_{m},\bm{\sigma}_{m}\in\mathds{R}^{C}$ are the batch statistics, and $\bm{\beta}_{m},\bm{\gamma}_{m}\in\mathds{R}^{C}$ are the affine parameters. The arithmetic is computed element-wise. In fact, such preference for invertible and spatially decomposable decoders may promote learning of the true disentangled representations [10, 88]. An illustration of the encoder and decoder architecture can be found in Section D of the Supplementary Material.
Finally, the expectation $\mathbb{E}_{q(\bm{z},\bm{v}\mid\bm{u};\bm{\psi})}[\log p(\bm{u}\mid\bm{z},\bm{v};{\bm{\theta}})]$ is approximated by Monte-Carlo ancestral sampling, where the likelihood is modelled by a Laplace distribution, i.e.,
$p(\bm{u}\mid\bm{z},\bm{v};\bm{\theta})$ $\propto\prod_{j=1}^{N}\prod_{\bm{\omega}\in\Omega_{j}}\exp\left\{-\frac{\absolutevalue{u_{j}(\bm{\omega})-f(\bm{z};\bm{\theta}_{j})\circ\phi_{j}^{-1}(\bm{\omega})}}{b}\right\},$ (21)
@@
The proposed framework also encourages learning of disentangled representations formally defined by Higgins et al. [8], which formulates disentangled representations as a reflection on the decomposition of certain symmetry groups acting on the world states (or observations). This explains why we choose to decompose the latent variables of our model into the common anatomy and the corresponding spatial diffeomorphisms: because they reflect two separate symmetry structures of the observed images, i.e., the ontological transformation changing the underlying anatomy, and the diffeomorphic transformation characterising geometric variation. A more thorough discussion of disentangled representations and their identifiability can be found in Section E of the Supplementary Material.
@@ 4.4 Learning, Inference and Scalability
Learning of our model is performed in an end-to-end fashion. All model parameters, including the auto-encoders and the registration modules, are trained simultaneously by stochastic gradient ascent to maximise the ELBO. In the inference stage, the transformations that register the image group are given by $\widehat{\bm{\phi}}=\{\widehat{\phi}_{j}\}_{j=1}^{N}$, where $\widehat{\phi}_{j}=\exp(\widehat{\bm{\mu}}_{j}^{+})$ and $\widehat{\bm{\mu}}_{j}^{+}$ is aggregated from the means of the velocity distributions at all levels predicted by the registration modules.
Remarkably, the proposed disentanglement of anatomy and geometry endows our model with powerful scalability. Since geometric and arithmetic averaging can be performed on any number (greater than 1) of representations, our model is capable of processing image groups of arbitrary sizes during training and testing. Let the number of images in an input group be $N^{\prime}=\sum_{m=1}^{M}{N_{m}}$, where $N_{m}$ is the number of images of modality $m$. Each image of modality $m$ is processed by the encoder for the corresponding modality to extract structural representations, and then the geometric and arithmetic means are calculated on the representations of all images to estimate the KL divergence and predict deformations for each image. For decoding, the reconstruction from the decoder of modality $m$ is inversely warped for $N_{m}$ times by the deformations for the $N_{m}$ images of modality $m$ to reconstruct original images. Thus, our model allows $N_{m}$’s to be different or even 0, and only requires $N^{\prime}\geq 2$ (for calculating averages) and at least two modalities are available in each group, i.e., $\sum_{m}\text{1}(N_{m}>0)\geq 2$ (for effectively learning modality-invariant representations).
@@
  • During evaluation, our model can register groups with arbitrary sizes, enabling large-scale and variable-size multi-modal groupwise registration, as explored in Section 5.6. In contrast, conventional iterative methods optimise similarity metrics with computational complexities that escalate rapidly with group size. Deep learning baselines, on the other hand, often have fixed channel numbers to take as input the concatenation of images [89], necessitating groups to be of uniform size. These limitations significantly restrict their applicability.
@@ 5 Experiments and Results
We evaluated our framework on a total of four publicly available datasets, including BraTS-2021 [90, 91, 92], MS-CMRSeg [93], Learn2Reg Abdominal MR-CT [94, 95, 96], and the OASIS dataset [97]. The experiments were implemented in PyTorch [98] and conducted on $\text{NVIDIA}^{\circledR}$ $\text{RTX}^{\textnormal{TM}}$ 3090 GPUs. The experimental materials, baselines, evaluation metrics, and results are detailed as follows.
@@ 5.2 Compared Methods
We compared two types of baselines with our models in the experiments. The first type features state-of-the-art iterative methods for multi-modal groupwise registration, including accumulated pairwise estimates (APE) [4], conditional template entropy (CTE) [5], and $\mathcal{X}$-CoReg [6]. These methods do not need training, but they can be prone to image artefacts and can hardly register large image groups due to excessive computational burdens. The second type of baselines extends the first type, aiming to learn the groupwise spatial correspondence via a neural network using the aforementioned similarity measures. To promote a fair comparison, we used Attention U-Net [83] for the network backbone, the same architecture as the encoder in our proposed framework. However, note that these learning-based baselines can only be trained to register images of a fixed group size, and therefore cannot be applied to data of variable group sizes in the test stage.
@@
  • Ours-CN: with complete learning and convolutional network-based registration modules.
  • Ours-PD: with partial learning and Demons-based registration modules.
  • Ours-CD: with complete learning and Demons-based registration modules.
Here, “C” and “P” signify Complete and Partial learning strategies, respectively; “N” and “D” indicate Network- or Demons-based registration modules, respectively.
@@ 5.3 Implementation Details
During preprocessing, the intensity range of each image was linearly normalised to $[0,1]$. For cardiac and abdominal images, they were further multiplied by the ROI mask to confine the information used by the network to predict registration. The encoder has an Attention U-Net architecture with $L=5$ levels, and the $l$-th level consists of two Conv-BN-LeakyReLU blocks with kernel size 3 and $16\times 2^{L-l}$ channels. The output of each convolutional block in the upsampling path is passed to a Conv layer to produce logits of the categorical structural distribution, with $8\times 2^{L-l}$ channels. Samples of the geometric mean structural distribution are passed to the decoder to reconstruct the original images. The decoder is composed of Conv-BN-LeakyReLU-Upsample blocks with kernel size 1 and $8\times 2^{L-l}$ channels.
To facilitate learning, different weights were used for reconstruction loss, intrinsic structural distance, and registration regularisation, corresponding to a likelihood-tempered ELBO [102]. The balancing weights of the loss terms for different training datasets are presented in the Supplementary Material Section G. The hyperparameter for spatial regularisation was set to $\lambda=10$ as in [46], and the maximum magnitude of the velocity calculated by the Demons algorithm was set to $\alpha_{0}^{l}=10\times 2^{l-L}$ at each level. The number of samples for the GR estimator was set to $S=3$ in all experiments. Network optimisation was performed by stochastic gradient descent using the Adam optimiser [103]. Besides, we used a learning rate of $1\times 10^{-3}$ and a batch size of 20 or 2 for the MS-CMRSeg or the other datasets, respectively.